I need to determine the line that passes through the points (3,9) and is parallel to the line 7 y - 28x = 2. The answer needs to be in slope intercept form

Any help would be great !!

Answer by fractalier(6550)

The slope of the line you want is equal to the slope of 7y - 28x = 2, since parallel lines have equal slopes. The slope here is found by solving for y:
7y = 28x = 2
y = 4x + 2/7
Thus m = 4
Now use the point-slope formula:
y - y1 = m(x - x1)
y - 9 = 4(x - 3)
y - 9 = 4x - 12
y = 4x - 3